OpenStudy (anonymous):

y^-6 z^0 -------- y^-3

OpenStudy (anonymous):

= y^(3)/y^(6), the negatives inverse their place in the fraction. any exponent raised to the 0 = 1 so, z^0 is out of the equation. simplifying the expression noted, = 1/y^3 = y^(-3). both of those are the same just written differently.